Find the distance between the following pair of points
A) (-1,0) and (3,0)
B) (1,-2) and (1,4)
C) (-2,3) and (2,0)
E) (a,a) and (-a,-a)
F) (a,b) and (-a,-b)
1
Expert's answer
2020-08-24T18:42:42-0400

In all cases we use the formula for the distance "d[(x_1,y_1),(x_2,y_2)]=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}".

A) "d=\sqrt{(3-(-1))^2+(0-0)^2}=\sqrt{4^2}=4."

B) "d=\sqrt{(1-1)^2+(4-(-2))^2}=\sqrt{6^2}=6."

C) "d=\sqrt{(2-(-2))^2+(0-3)^2}=\sqrt{25}=5."

E) "d=\sqrt{(-a-a)^2+(-a-a)^2}=\sqrt{8a^2}=2\sqrt{2}a."

F)

"d=\sqrt{(-a-a)^2+(-b-b)^2}=\sqrt{4a^2+4b^2}=\\
2\sqrt{a^2+b^2}."